bshafer:

Xavier is a salesperson who is paid a fixed amount of $455 per week. He also earns a commission of 3% on the sales he makes. If Xavier wants to earn more than $575 in one week, how many dollars (x) in sales must he make? A. x > 1,060 B. x > 4,000 C. x < 3,600 D. x < 1,060

umm:

He would want to earn more than $575. $575 = $455 + $120 so he'll need to earn at least $120. Commission = 3% on sales. Let the sales be $x Commission = 3% of $x \[120 = \frac{ 3 }{ 100 }\times x\] \[\frac{ 120 \times 100}{ 3 }= x \] \[x=4,000\]